ws (Indirect Method) The Wolf Companys income statement of Cash Flows ve balance sheets at Dec nt ember 31 of 2016 and 2015 are shown below: income statement and com LO2,3 WOLFF COMPANY Income Statement the Year Ended December 31,2016 Sales revenue Cost of goods sold Wages expense Insurance expense. 86,000 12,000 13,000 12,000 ncome tax expense Net income. 29,000 582,000 63,000 WOLFF COMPANY Balance Sheets Dec. 31, Dec. 31, 2015 s 52,000 8,.000 Accounts receivable Prepaid insurance?.. , Plant assets 90,000 60,000 7,000 219,000 195,000 .. $339,000 $247,000 Total assets Liabilities and Stockholders Equity Wages payable Bonds payable Retained eanings.. Total liabilities and stockholders equity 7000 $ 10.000 6,000 7,000 000 75,000 9,000 6,000 90,000 90,000 59,000 $339,000 $247.000 . . .. nt assets were purchased for cash and bonds payable were issued for cash. Bond interest is paid semi-annually on June 30 and December 31. Cash dividend s of S36,000 were declared and paid during 2016. Pla Accounts payable relate to merchandise purchases. Required d Calculate the change in cash that occurred during 2016 c. Compute free cash flow Prepare a statement of cash flows using the indirect method. compute the operating-cash-flow-to-current-liabilities ratio. compute the operating-cash-flow-to-capital-expenditures ratio
Answer #1

PART-1)

Change in Cash during 2016 $44,000 increase

PART-2)

Cash Flow Provided by Operating Activities

Working

Net income

63,000

Add/Minus

Depreciation

13,000

Accounts Receivable

Increase

-9,000

32,000 - 41,000

Inventory

Increase

-30,000

30,000 - 60,000

Prepaid Insurance

Decrease

2,000

7,000 - 5,000

Accounts Payable

Decrease

-3,000

7,000 - 10,000

Wage expense

Increase

3,000

9,000 - 6,000

Income Tax Payable

Decrease

-1,000

6,000 - 7,000

Cash Flow Provided by Operating Activities

38,000

Cash Flow from Investing Activities

Purchase of Plant Assets

24,000

195000-219000=-24000

Cash Flow from Financing Activities

Issuance of Bonds Payable

66,000

141,000-75,000=66000

Payment of Dividends

-36,000

Cash Flow from Financing Activities

30,000

Net increase in cash

44,000

38,000-24,000+30,000

Cash at Beginning of Year

8,000

Cash at End of Year

52,000

PART-3)

Free cash flow $ -14,000

Working: Operating cash flow - Purchase of asset = 38,000 - 24,000 = 14,000

PART-4)

Operating-cash-flow-to-current-liabilities ratio : 1.69

Working:

operating cash flow/Average current liabilities

=38,000 / {(7000+10000+9000+6000+6000+7000)/2}

=38,000 / 22,500

=1.69

PART-5)

Operating-cash-flow-to-capital-expenditures ratio : 1.58

Working:

Operating-cash-flow / Capital-expenditures = 38,000 / 24,000 = 1.58